OpenStudy (anonymous):

Please help! Verify the identity. 4 csc 2x = 2 csc2x tan x

OpenStudy (anonymous):

@Nnesha @jango_IN_DTOWN ???

OpenStudy (anonymous):

Can you help me solve this??

OpenStudy (jango_in_dtown):

wait a bit

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (jango_in_dtown):

is it \[4\csc 2x=2\csc ^2xtanx \]

OpenStudy (jango_in_dtown):

\[4\csc 2x=\frac{ 4 }{ \sin 2x}=\frac{ 4 }{ 2\sin x \cos x }\]

OpenStudy (anonymous):

\[4 \csc 2x = 2 \csc^2x \tan x\]

OpenStudy (jango_in_dtown):

\[=2\csc x\frac{ \sin x }{\cos x }\frac{ 1 }{ \sin x }\]

OpenStudy (jango_in_dtown):

\[=2\csc^2xtan x\]

OpenStudy (anonymous):

Can you explain how you got that answer.

OpenStudy (jango_in_dtown):

I multiplies numerator and denominator by sin x and 1/sin x =csc x and sin x/cos x=tan x

OpenStudy (jango_in_dtown):

*multiplied

OpenStudy (jango_in_dtown):

sin 2x=2 sin x cos x

OpenStudy (anonymous):

Thanks a million again!! @jango_IN_DTOWN

OpenStudy (jango_in_dtown):

write the reasons wherever you use other identities.. The teacher will be happy too.:)